Let G be the set of all 3×3 non-singular matrices with entries from Q. Let X be a fixed matrix in G.prove that operation *defined by A*B= X^(-1)ABX be a binary operation. How do you prove that G is a group with respect to binary operation defined by A*B=X^(-1) ABX?
1
Expert's answer
2020-09-28T19:29:51-0400

"G=\{A=\begin{pmatrix}
 a_{11} & a_{12}&a_{13} \\
 a_{21} & a_{22}&a_{23} \\
 a_{31} & a_{32}&a_{33} 
\end{pmatrix}, a_{ij}\in Q,\\
 i=1,2,3,j=1,2,3, detA\neq0 \},\\
X=\begin{pmatrix}
 x_{11} & x_{12}&x_{13} \\
 x_{21} & x_{22}&x_{23} \\
 x_{31} & x_{32}&x_{33}
\end{pmatrix}, detX\neq0"

X-fixed matrix in G

"detX\neq0\to\exists X^{-1}, detX^{-1}\neq0"

"X^{-1}=\frac{1}{detX}\begin{pmatrix}
 y_{11} & y_{12}&y_{13} \\
 y_{21} & y_{22}&y_{23} \\
 y_{31} & y_{32}&y_{33}
\end{pmatrix}"

where "y_{ij}" - algebraic addition to the element "x_{ij}" .

"B\in G: B=\begin{pmatrix}
 b_{11} & b_{12}&b_{13} \\
 b_{21} & b_{22}&b_{23} \\
 b_{31} & b_{32}&b_{33} 
\end{pmatrix}, b_{ij}\in Q,\\
 detB\neq0"

"A*B=X^{-1}ABX"

"X^{-1},A,B,X" 3 "\times" 3 matrices

"A*B" 3"\times" 3 matrices with elemets in Q

"det(A*B)=det(X^{-1}ABX)=\\
=det(X^{-1})det(A)det(B)det(X)\neq0\\
A*B\in G"

So "*" a binary operation.

Prove that G is a group.

1. "A*B\in G"

2."(A*B)*C=A*(B*C)"

3."I=\begin{pmatrix}
 1& 0&0 \\
 0& 1&0\\
0&0&1
\end{pmatrix}\in G,1\in Q, det(I)=1\neq0"

4."A*A^{-1}=I"

Find "A^{-1}:"

"A*A^{-1}=X^{-1}AA^{-1}X=X^{-1}IX=X^{-1}X=I\\
det(A^{-1})\neq0"

So G is a group with respect to binary operation